### v1

@@ Abstract
Reinforcement learning (RL) is a powerful tool for decision-making in uncertain environments, but it often requires large amounts of data to learn an optimal policy. We propose using prior model knowledge to guide the exploration process to speed up this learning process. This model knowledge comes in the form of a model set to which the true transition kernel and reward function belong. We optimize over this model set to obtain upper and lower bounds on the Q-function, which are then used to guide the exploration of the agent. We provide theoretical guarantees on the convergence of the Q-function to the optimal Q-function under the proposed class of exploring policies. Furthermore, we also introduce a data-driven regularized version of the model set optimization problem that ensures the convergence of the class of exploring policies to the optimal policy. Lastly, we show that when the model set has a specific structure, namely the bounded-parameter MDP (BMDP) framework, the regularized model set optimization problem becomes convex and simple to implement. In this setting, we also show that we obtain finite-time convergence to the optimal policy under additional assumptions. We demonstrate the effectiveness of the proposed exploration strategy in a simulation study. The results indicate that the proposed method can significantly speed up the learning process in reinforcement learning.
@@ I INTRODUCTION
Reinforcement learning has shown great success in solving complex decision-making problems in various domains, such as robotics, games, and finance [1]. However, RL often requires large amounts of data to learn an optimal policy, which can be impractical in many real-world applications. This is especially true in environments where data collection is expensive or time-consuming, such as in robotics or healthcare. The need for large datasets comes from the fact that RL algorithms typically explore the environment randomly, causing the agent to visit many suboptimal states and actions [2].
@@
A key opportunity could arise when having access to model knowledge, which is often the case in many scenarios. This information is typically not used in traditional RL algorithms. For instance, we may know that particular states cannot be reached from certain other states or that there is a specific goal state. In this paper, we propose using such available knowledge, giving information on the set of models that the true model belongs to, to guide the exploration of the agent, thus speeding up the learning process.
@@
More specifically, we propose to estimate bounds on the Q-function using optimistic/pessimistic optimization over the available model set. We then use these bounds to guide the agent’s exploration by prioritizing uncertain or promising regions of the state/action space. Additionally, we propose regularizing the model set optimization problem using a database of observed transitions. We show that the Q-function converges to the optimal Q-function under the exploring policy and that the exploring policy converges to the optimal policy in the regularized case. Although these results hold in the general case, we show that for specific model set characterizations, we obtain a practical algorithm with a convex (regularized) model set optimization.
@@
This work is at the intersection of several research areas in reinforcement learning. We will now briefly review the literature on robust reinforcement learning, smart exploration in reinforcement learning, and model-based reinforcement learning, making connections to the content of the present paper.
In robust reinforcement learning, we typically aim to find policies that are robust to model uncertainty. This can be done by optimizing over a set of models, as in [3, 4, 5, 6]. These works typically focus on adversarial optimization over the model set to find a policy that is robust to the worst-case model in the set. Our work differs in that we use the model set to guide the exploration of the agent, rather than to find a robust policy.
Smart exploration in reinforcement learning is a well-studied topic, with many works focusing on how to explore the environment efficiently to speed up the learning process. Some works focus on using uncertainty estimates to guide the exploration, such as [2, 7, 8, 9]. Our work differs in that we use a set of models to guide the exploration, rather than using uncertainty estimates.
Model-based reinforcement learning is a popular approach to speed up the learning process in RL. In model-based RL, the agent learns a model of the environment and uses this model to plan its actions [10, 11]. Our work differs in that we use the model knowledge to guide the exploration of the agent, rather than to plan its actions.
@@
This work also relates to two key ideas in the literature. The first is the bounded-parameter MDP (BMDP) framework [12] (also known as uMDPs or IMDPs) which characterizes the uncertainty in the model set through intervals. The BMDP framework has been used to develop algorithms that are robust to model uncertainty, such as [6, 13, 14]. The present work builds on this idea by using the BMDP framework to guide the exploration of the agent. The second is Bayesian optimization, which is a popular method for optimizing (static) black-box functions [15] by balancing exploration and exploitation in a Bayesian framework. Bayesian RL builds on this by using Bayesian inference to guide the agent’s exploration [16]. Our work uses similar heuristics to guide the exploration of the agent but differs in that we also use model knowledge.
The main contributions of this paper can be summarized as follows.
@@
  • We introduce a novel exploration strategy for reinforcement learning that leverages bounds on the Q-function obtained from a model set. We provide general theoretical guarantees on the convergence of the Q-function to the optimal Q-function under the proposed exploration strategy.
@@
  • We propose a practical algorithm based on the BMDP framework and show that, under specific conditions, we obtain finite-time convergence to the optimal policy.
  • We demonstrate the effectiveness of our approach in a numerical case study.
The remainder of this paper is organized as follows. In Section II, we introduce the problem formulation. In Section III, we present the general framework for the proposed method and detail most of the technical results. Section IV discusses a practical algorithm that makes use of the theoretical results. Section V discusses a brief simulation study and Section VI concludes the paper.
@@ II PROBLEM FORMULATION
  • Action Space $\mathcal{U}$ is a measurable (Polish) space representing the set of possible inputs or actions that the agent can take. We denote the Borel $\sigma$-algebra on $\mathcal{U}$ by $\mathcal{B}(\mathcal{U})$.
@@
  • Reward Function $g:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is a bounded reward function, where $g(x,u)$ is the reward obtained by the agent when taking action $u$ in state $x$.
@@
The goal in this setting is to find a policy $\pi\in\Pi$, where $\Pi:=\{\pi:\mathcal{X}\to\text{Prob}(\mathcal{U})\mid\pi\text{ is measurable}\}$, such that the expected cumulative reward or value is maximized. The value function $V^{\pi}:\mathcal{X}\to\mathbb{R}$ of a policy $\pi\in\Pi$ is defined as
@@
where $Q^{\pi}(x,u)$ is the expected cumulative reward of taking action $u$ in state $x$, then following policy $\pi$. We assume that for every $\pi\in\Pi$, the integrals in (1) and in (2) are well-defined. The optimal Q-function $Q^{*}:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is defined as
@@
An optimal policy $\pi^{*}$ can be obtained from the optimal Q-function as $\pi^{*}(x)\in\arg\max_{u\in\mathcal{U}}Q^{*}(x,u)$. The optimal Q-function can also be defined in terms of the Bellman optimality operator $\mathcal{T}$ as $Q^{*}=\mathcal{T}Q^{*}$, where the Bellman operator $\mathcal{T}:\mathcal{B}(\mathcal{X}\times\mathcal{U})\to\mathcal{B}(\mathcal{X}\times\mathcal{U})$ is defined as
$(\mathcal{T}Q)(x,u)=g(x,u)+\gamma\int_{\mathcal{X}}\sup_{u^{\prime}\in\mathcal{U}}Q(x^{\prime},u^{\prime})P(x,u,dx^{\prime}).$ (4)
In reinforcement learning, our objective is to find the optimal policy $\pi^{*}$ by interacting with the environment and learning from the observed data $\mathcal{D}:={\{}(x_{k},u_{k},x_{k+1}){\}}_{k=0}^{N-1}$. The transition kernel $P$ and the reward function $g$ are assumed to be unknown to the agent. However, we assume that the agent has access to a model set $\hat{\mathcal{M}}$ that the true transition kernel and reward function belong to. This is formalized in the following assumption.
@@
The model set $\hat{\mathcal{M}}$ is defined as ^$\hat{\mathcal{M}}=(\mathcal{X},\mathcal{U},\mathcal{P},\mathcal{G},\gamma)$, where $\mathcal{P}$ is a set of transition kernels and $\mathcal{G}$ is a set of reward functions. The true transition kernel $P$ belongs to $\mathcal{P}$, and the true reward function $g$ belongs to $\mathcal{G}$. The model set $\hat{\mathcal{M}}$ is known to the agent.
@@ III PROPOSED METHOD
We will first discuss a version of the proposed method that fits the general MDP setting introduced in the previous section. This version cannot be directly applied in practice because an optimization over the general model set is intractable. This version presents the theoretical foundation for a more practical version that we will detail in Section IV.
@@ III-A General Framework
Let $\underaccent{\bar}{\mathcal{T}}$ and $\bar{\mathcal{T}}$ denote the Bellman update operators defined in (5). These operators are contraction mappings on $\mathcal{B}(\mathcal{X}\times\mathcal{U})$. Let $\underaccent{\bar}{Q}^{*}$ and $\bar{Q}^{*}$ be the unique fixed points of $\underaccent{\bar}{\mathcal{T}}$ and $\bar{\mathcal{T}}$, respectively, obtained by repeated application of these operators. Then, the fixed points satisfy
@@
where $Q^{*}$ is the optimal Q-function defined in (3). Consequently, if we define the corresponding value functions as
$\underaccent{\bar}{V}^{*}(x):=\sup_{u\in\mathcal{U}}\underaccent{\bar}{Q}^{*}(x,u)\quad\text{and}\quad\bar{V}^{*}(x):=\sup_{u\in\mathcal{U}}\bar{Q}^{*}(x,u),$ (8)
then we have
$\underaccent{\bar}{V}^{*}(x)\leq V^{*}(x)\leq\bar{V}^{*}(x),\quad\forall x\in\mathcal{X},$ (9)
where $V^{*}(x):=\max_{u\in\mathcal{U}}Q^{*}(x,u)$ is the optimal value function.
@@
Since the reward function $g$ is bounded and $\gamma\in[0,1)$, the space $\mathcal{B}(\mathcal{X}\times\mathcal{U})$ of bounded measurable functions, equipped with the supremum norm, is a Banach space. As shown in [17], the standard Bellman operator defined in (4) is a contraction mapping with modulus $\gamma$, i.e.,
$\|\mathcal{T}Q_{1}-\mathcal{T}Q_{2}\|_{\infty}\leq\gamma\|Q_{1}-Q_{2}\|_{\infty},$ (10)
for any bounded functions $Q_{1}$ and $Q_{2}$. Using similar reasoning, the modified operators $\underaccent{\bar}{\mathcal{T}}$ and $\bar{\mathcal{T}}$—which differ from $\mathcal{T}$ only by taking the infimum and supremum over the model set, respectively—are also contractions with the same modulus $\gamma$. Hence, by the Banach fixed-point theorem, each operator has a unique fixed point, namely $\underaccent{\bar}{Q}^{*}$ for $\underaccent{\bar}{\mathcal{T}}$ and $\bar{Q}^{*}$ for $\bar{\mathcal{T}}$.
Since the true MDP $\mathcal{M}$ is contained in the model set $\hat{\mathcal{M}}$ (with $g\in\mathcal{G}$ and $P\in\mathcal{P}$), for any bounded Q-function $Q$ we have
$\underaccent{\bar}{\mathcal{T}}Q(x,u)\leq\mathcal{T}Q(x,u)\leq\bar{\mathcal{T}}Q(x,u),\quad\forall(x,u)\in\mathcal{X}\times\mathcal{U}.$ (11)
The monotonicity of these operators then implies that their unique fixed points are ordered as in (7), where $Q^{*}$ is the fixed point of $\mathcal{T}$, i.e., the optimal Q-function. The value function bounds (9) follow directly from the Q-function bounds (7) and their definition (8).
This completes the proof. ∎
The paper [12] on bounded-parameter MDPs establishes similar guaranteed (optimistic and pessimistic) bounds for the value function in the finite-state case for a particular model set structure. Our result generalizes this idea by considering the Q-function and extending the setting to infinite state spaces. The key is that even in the infinite-state scenario, provided that the boundedness and contraction conditions hold (as detailed in [17]), the same sandwiching property can be proven.
Note that based on the guaranteed Q-bounds, we can in some cases guarantee that certain inputs are suboptimal or optimal. This is formalized in the following proposition.
@@
Let $\underaccent{\bar}{Q}^{*}$ and $\bar{Q}^{*}$ be guaranteed lower and upper bounds on the optimal Q-function $Q^{*}$, respectively, such that (7) holds. Then, the input $u$ at state $x$ is guaranteed to be suboptimal if
@@
and the input $u$ at state $x$ is guaranteed to be optimal if
@@
Proof.
For suboptimality, note that $\bar{Q}^{*}(x,u)$ is the best-case return of action $u$, while $\underaccent{\bar}{V}^{*}(x)$ is the best worst-case return across actions at $x$. Hence, if (12) holds, even the most favorable outcome for $u$ is inferior to the least favorable outcome of at least one other action, ensuring that $u$ is suboptimal.
For optimality, if (13) holds, then the worst-case return for $u$ is at least as high as the best-case return for every other action. Since $Q^{*}(x,u)$ lies between these bounds, $u$ must yield the highest possible Q-value at $x$ and is therefore optimal. ∎
@@ III-B Exploration Strategy
$\phi(x,u)=\begin{cases}\xi,&\text{if }\underaccent{\bar}{Q}(x,u)\geq\max_{v\in\mathcal{U}\setminus\{u\}}\bar{Q}(x,v)\\ \beta(x,u),&\text{if }\underaccent{\bar}{Q}(x,u)\neq\bar{Q}(x,u)\text{ and }\bar{Q}(x,u)>\underaccent{\bar}{V}(x)\\ 0,&\text{if }\underaccent{\bar}{Q}(x,u)=\bar{Q}(x,u)\text{ and }\bar{Q}(x,u)\leq\underaccent{\bar}{V}(x)\\ \zeta,&\text{otherwise},\end{cases}$ (14)
with $\phi:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{\geq 0}$ the sampling weight of input $u$ at state $x$. Here, $\xi\in\mathbb{R}_{>0}$ and $\zeta\in\mathbb{R}_{\geq 0}$ are constants, and $\beta:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{>0}$. The induced exploration policy $\pi^{\phi}$ is defined by
@@
provided that $\int_{\mathcal{U}}\phi(x,v)\,\mu(dv)>0$, with $\mu$ a suitable reference measure on $\mathcal{U}$ (e.g., the counting measure if $\mathcal{U}$ is finite).
The first case in (14) implies that the input $u$ at state $x$ is guaranteed to be optimal (by Proposition 3). For this reason, we assign it the positive weight $\xi$. Note that if there exists an input that satisfies this condition, this rules out the possibility that any other inputs satisfy the second condition in (14). The second case implies that the quality of input $u$ at state $x$ is uncertain but that the input has the potential to improve $\underaccent{\bar}{V}$. We assign this choice of input a positive weight $\beta$ that depends on $x$ and $u$. This allows us to potentially assign more weight to the input if it is highly uncertain or has a high Q-value upper bound $\bar{Q}$. Section III-B1 provides heuristic guidance on the choice of $\beta$, inspired by Bayesian optimization. The third case in (14) assigns zero weight to actions with tight Q-bounds that are guaranteed to be suboptimal, while the fourth case assigns a nonnegative weight of $\zeta$ to inputs which are uncertain yet appear to be suboptimal. Note that in many cases, we can select $\zeta=0$ to prune such inputs. However, this pruning might be too aggressive in some cases if we introduce data-driven regularization of the model set optimization, which will be discussed later in this paper. An illustration representing the exploring policy based on the bounds is provided in Fig. 1.
@@
The convergence properties of standard Q-learning are retained by the proposed exploring policy, as the exploration is based on guaranteed bounds. To show this, we provide the following Theorem.
@@
Let $Q$ be the Q-function that is obtained using standard Q-learning, i.e.,
@@
under the data-gathering policy given by the exploring policy $\pi^{\phi}$ defined in (15) with arbitrary parameters $\xi\in\mathbb{R}_{>0},\zeta\in\mathbb{R}_{\geq 0}$ and function $\beta:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{>0}$. If every state is visited infinitely often (using, e.g., exploring starts) and the learning rate $\alpha_{k}$ satisfies the Robbins-Monro conditions, i.e., $\sum_{k=0}^{\infty}\alpha_{k}=\infty$ and $\sum_{k=0}^{\infty}\alpha_{k}^{2}<\infty$, then the greedy policy with respect to $Q$ converges to the optimal policy with probability 1 as $k\to\infty$
@@
Standard Q-learning convergence (e.g., [18, 17]) guarantees that if every state–action pair is updated infinitely often, then $Q$ converges to $Q^{*}$. In our setting, by design, the exploring policy $\pi^{\phi}$ prunes actions that are guaranteed to be suboptimal by (12). Although this means that not every state–action pair is updated infinitely often, it still means that every state $x$, every action $u$ that could be optimal is updated infinitely often (via exploring starts).
Thus, even if $Q(x,u)$ does not converge to $Q^{*}(x,u)$ for all $(x,u)$, the greedy policy $\pi(x)\in\arg\max_{u\in\mathcal{U}}Q(x,u)$ converges to optimal policy $\pi^{*}(x)$ with probability 1, since for every state $x$, the Q-value of every candidate action either converges to its true optimal Q-value or is strictly dominated by another action that does. ∎
@@
III-B1 Choice of Weighting Function $\beta$

### v2

@@ Abstract
Reinforcement learning (RL) is a powerful framework for decision-making in uncertain environments, but it often requires large amounts of data to learn an optimal policy. We address this challenge by incorporating prior model knowledge to guide exploration and accelerate the learning process. Specifically, we assume access to a model set that contains the true transition kernel and reward function. We optimize over this model set to obtain upper and lower bounds on the Q-function, which are then used to guide the exploration of the agent. We provide theoretical guarantees on the convergence of the Q-function to the optimal Q-function under the proposed class of exploring policies. Furthermore, we also introduce a data-driven regularized version of the model set optimization problem that ensures the convergence of the class of exploring policies to the optimal policy. Lastly, we show that when the model set has a specific structure, namely the bounded-parameter MDP (BMDP) framework, the regularized model set optimization problem becomes convex and simple to implement. In this setting, we also prove finite-time convergence to the optimal policy under mild assumptions. We demonstrate the effectiveness of the proposed exploration strategy, which we call BUMEX (Bounded Uncertainty Model-based Exploration), in a simulation study. The results indicate that the proposed method can significantly accelerate learning in benchmark examples. A toolbox is available at https://github.com/JvHulst/BUMEX.
@@ I INTRODUCTION
Reinforcement learning (RL) has shown great success in solving complex decision-making problems in various domains, such as robotics, games, and finance [1]. However, RL often requires large amounts of data to learn an optimal policy, which can be impractical in many real-world applications. This is especially true in environments where data collection is expensive or time-consuming, such as in robotics or healthcare. The need for large datasets comes from the fact that RL algorithms typically explore the environment randomly, causing the agent to visit many suboptimal states and actions [2].
@@
A key opportunity arises when one has access to model knowledge, which is often the case in many scenarios. This information is typically not utilized in traditional RL algorithms. For instance, we may know that certain states cannot be reached from particular other states or that there is a specific goal state. In this paper, we propose using such available knowledge, giving information on the set of models to which the true model belongs, and using this to guide the agent’s exploration, thus speeding up the learning process.
@@
More specifically, we propose estimating bounds on the Q-function using optimistic/pessimistic optimization over the available model set. We then use these bounds to guide the agent’s exploration by prioritizing uncertain or promising regions of the state/action space. Additionally, we propose regularizing the model set optimization problem using a database of observed transitions. We show that the Q-function converges to the optimal Q-function under the exploring policy and that the exploring policy converges to the optimal policy in the regularized case. Moreover, for a specific model set characterization we obtain a practical algorithm with a convex (regularized) model set optimization. We refer to this practical algorithm as BUMEX (Bounded Uncertainty Model-based Exploration).
@@
This work is at the intersection of model-based RL and smart exploration in RL. We will now briefly review the relevant literature, highlighting the relation and differences to the current work.
Model-based reinforcement learning is a popular approach to speed up the learning process in RL by incorporating model knowledge. In traditional model-based RL, the agent learns a model of the environment and uses this model to plan its actions [3, 4]. In our work, we use prior model knowledge to guide the exploration of the agent, rather than to plan its actions. Smart exploration in RL focuses on how to explore the environment efficiently to speed up learning. Several prominent approaches use confidence bounds on value functions to guide exploration, including Upper Confidence Bound methods [2, 5], UCRL2 [6], and Posterior Sampling for Reinforcement Learning (PSRL) [7]. Bayesian RL builds on this by using Bayesian inference to guide the agent’s exploration [8]. These methods typically rely on uncertainty estimates derived from a single nominal model or confidence intervals around model parameters. In contrast, our work employs a set of models to direct exploration, rather than relying on a nominal model with uncertainty estimates.
@@
Our work builds most directly on the bounded-parameter MDP (BMDP) framework [9], which characterizes model uncertainty through intervals on transition probabilities and rewards. BMDPs are closely related to uncertain MDPs (uMDPs) and interval MDPs (IMDPs). The BMDP framework has been used to develop algorithms that are robust to model uncertainty [10]. The present work builds upon this idea by using the BMDP framework to guide the exploration of the agent. Additionally, our exploration strategy draws inspiration from Bayesian optimization [11], which balances exploration and exploitation by using acquisition functions. We adapt similar heuristics for selecting actions based on Q-function bounds.
The main contributions can be summarized as follows.
@@
  • We introduce an exploration strategy for reinforcement learning that leverages bounds on the Q-function obtained from a model set. We provide theoretical guarantees on the convergence of the Q-function to the optimal Q-function under the proposed exploration strategy.
@@
  • We propose a practical algorithm, BUMEX, for the BMDP framework and establish, under specific conditions, finite-time convergence to the optimal policy.
Our main theoretical results focus on finite state-action spaces, which allows us to provide (finite-time) convergence guarantees and construct a practical algorithm.
The paper is organized as follows. Section II, introduces the problem formulation. Section III presents the general framework for the proposed method and the technical results. Section IV discusses a practical algorithm that leverages on the theoretical results. Section V presents simulations and Section VI concludes the paper.
@@ II PROBLEM FORMULATION
  • Action Space $\mathcal{U}$ is a measurable (Polish) space representing the set of inputs or actions that the agent can take. We denote the Borel $\sigma$-algebra on $\mathcal{U}$ by $\mathcal{B}(\mathcal{U})$.
@@
  • Reward Function $g:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is a bounded, measurable function, where $g(x,u)$ is the reward obtained by the agent when taking action $u$ in state $x$. We assume deterministic rewards for simplicity. The extension to stochastic rewards is straightforward.
@@
The goal in this setting is to find a policy $\pi\in\Pi$, where $\Pi:=\{\pi:\mathcal{X}\to\text{Prob}(\mathcal{U})\mid\pi\text{ is measurable}\}$, such that the expected cumulative reward or value is maximized. Here $\text{Prob}(\mathcal{U})$ denotes the set of Borel probability measures on $(\mathcal{U},\mathcal{B}(\mathcal{U}))$. The value function $V^{\pi}:\mathcal{X}\to\mathbb{R}$ of a policy $\pi\in\Pi$ is defined as
@@
where $Q^{\pi}(x,u)$ is the expected cumulative reward of taking action $u$ in state $x$, then following policy $\pi$. We assume that for every $\pi\in\Pi$, the integrals in (1), (2) are well-defined. The optimal Q-function $Q^{*}:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is defined as
@@
An optimal policy $\pi^{*}$ can be obtained from the optimal Q-function as $\pi^{*}(x)\in\arg\max_{u\in\mathcal{U}}Q^{*}(x,u)$. The optimal Q-function can also be defined in terms of the Bellman operator $\mathcal{T}$ as $Q^{*}=\mathcal{T}Q^{*}$, where $\mathcal{T}:\mathcal{B}(\mathcal{X}\times\mathcal{U})\to\mathcal{B}(\mathcal{X}\times\mathcal{U})$ is defined as
$(\mathcal{T}Q)(x,u)=g(x,u)+\gamma\int_{\mathcal{X}}\sup_{u^{\prime}\in\mathcal{U}}Q(x^{\prime},u^{\prime})P(x,u,dx^{\prime}).$
As usual in reinforcement learning, our objective is to find the optimal policy $\pi^{*}$ by interacting with the environment and learning from the observed data $\mathcal{D}:={\{}(x_{k},u_{k},x_{k+1}){\}}_{k=0}^{N-1}$. The transition kernel $P$ and the reward function $g$ are assumed to be unknown to the agent. However, we assume that the agent has access to a model set $\hat{\mathcal{M}}$ that contains the true transition kernel and reward function. This is formalized in the following assumption.
@@
The model set $\hat{\mathcal{M}}$ is defined as $\hat{\mathcal{M}}=(\mathcal{X},\mathcal{U},\mathcal{P},\mathcal{G},\gamma)$, where $\mathcal{P}$ is a set of transition kernels and $\mathcal{G}$ is a set of reward functions. The true transition kernel $P$ belongs to $\mathcal{P}$, and the true reward function $g$ belongs to $\mathcal{G}$. The model set $\hat{\mathcal{M}}$ is known to the agent.
@@ III PROPOSED METHOD
We will first discuss a version of the proposed method that fits the general MDP setting introduced in the previous section. This version cannot be directly applied in practice because optimization over the general model set is intractable. This version presents the theoretical foundation for a more practical version that we will detail in Section IV.
@@ III-A General Framework
Let $\underaccent{\bar}{\mathcal{T}}$ and $\bar{\mathcal{T}}$ denote the Bellman operators defined in (4). These operators are contraction mappings on $\mathcal{B}(\mathcal{X}\times\mathcal{U})$ with unique fixed points $\underaccent{\bar}{Q}^{*}$ and $\bar{Q}^{*}$, respectively. Then
@@
where $Q^{*}$ is the optimal Q-function defined in (3). Similarly, the corresponding value functions
$\underaccent{\bar}{V}^{*}(x):=\sup_{u\in\mathcal{U}}\underaccent{\bar}{Q}^{*}(x,u)\quad\text{and}\quad\bar{V}^{*}(x):=\sup_{u\in\mathcal{U}}\bar{Q}^{*}(x,u),$
satisfy $\underaccent{\bar}{V}^{*}(x)\leq V^{*}(x)\leq\bar{V}^{*}(x),\quad\forall x\in\mathcal{X}$, where $V^{*}(x):=\max_{u\in\mathcal{U}}Q^{*}(x,u)$ is the optimal value function.
@@
Since the reward function $g$ is bounded and $\gamma\in[0,1)$, the operators $\underaccent{\bar}{\mathcal{T}}$ and $\bar{\mathcal{T}}$ are contraction mappings with modulus $\gamma$ in finite spaces. By the Banach fixed-point theorem, each operator has a unique fixed point. Since the true MDP $\mathcal{M}$ is contained in the model set $\hat{\mathcal{M}}$, we have $\underaccent{\bar}{\mathcal{T}}Q\leq\mathcal{T}Q\leq\bar{\mathcal{T}}Q$ for any Q-function $Q$. The monotonicity of these operators implies the ordering of their fixed points as in (6), and the value function bounds follow directly. ∎
The paper [9] on bounded-parameter MDPs establishes similar bounds for the value function in finite state spaces using interval-based model sets. Our result generalizes this by considering both finite and infinite spaces and by using Q-function bounds instead. Note that the initialization of the Q-bounds does not affect the sandwiching property that is obtained upon convergence. Typically, we initialize the Q-bounds using prior knowledge or heuristics about the environment to speed up the convergence.
As stated next, the guaranteed Q-bounds can be used to test if certain inputs are suboptimal or optimal.
@@
Let $\underaccent{\bar}{Q}^{*}$ and $\bar{Q}^{*}$ be guaranteed lower and upper bounds on the optimal Q-function $Q^{*}$, respectively, i.e., (6) holds. Then, the input $u$ at state $x$ is suboptimal if
@@
and the input $u$ at state $x$ is optimal if
@@
This result follows directly from the guaranteed Q-bounds.
@@ III-B Exploration Strategy
[…]
with $\phi:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{\geq 0}$ the sampling weight of input $u$ at state $x$. Here, $\xi\in\mathbb{R}_{>0}$ and $\zeta\in\mathbb{R}_{\geq 0}$ are constants, and $\beta:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{>0}$. These conditions form a partition of the state-action space $\mathcal{X}\times\mathcal{U}$. The induced exploration policy $\pi^{\phi}$ is defined by
@@
with $\mu$ a suitable reference measure on $\mathcal{U}$ (e.g., the counting measure if $\mathcal{U}$ is finite).
The first case in (9) implies that the input $u$ at state $x$ is guaranteed to be optimal (by Proposition 3). For this reason, we assign it an arbitrary positive weight $\xi$. Note that if an input satisfies this condition, this rules out the possibility that any other inputs satisfy the second condition in (9).
The second case implies that the quality of input $u$ at state $x$ is uncertain but that the input has the potential to improve $\underaccent{\bar}{V}$. We assign this input a positive weight $\beta$ that depends on $x$ and $u$. This allows us to potentially assign more weight to the input if it is highly uncertain or has a high Q-value upper bound $\bar{Q}$. Section III-B1 provides heuristic guidance on the choice of $\beta$, inspired by Bayesian optimization.
The third case in (9) assigns zero weight to actions with tight Q-bounds that are guaranteed to be suboptimal. The fourth case assigns a nonnegative weight of $\zeta$ to inputs which are uncertain yet appear to be suboptimal, while the fifth case assigns the same weight to inputs which are certain and not guaranteed to be suboptimal. Note that in many cases, we can select $\zeta=0$ to prune such inputs. However, this pruning might be too aggressive in some cases if we introduce data-driven regularization of the model set optimization, which will be discussed later in this paper. Fig. 1 illustrates the exploring policy based on the bounds.
@@
As shown next, the convergence properties of standard Q-learning are retained by the proposed exploring policy, as the exploration is based on guaranteed bounds.
@@
Consider finite state-action spaces. Let $Q$ be the Q-function obtained using standard Q-learning
@@
under the exploring policy $\pi^{\phi}$ defined in (10) with parameters $\xi\in\mathbb{R}_{>0},\zeta\in\mathbb{R}_{\geq 0}$ and function $\beta:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{>0}$. If every state is visited infinitely often (through e.g., exploring starts), and the learning rate $\alpha_{k}$ satisfies the Robbins-Monro conditions $\sum_{k=0}^{\infty}\alpha_{k}=\infty$ and $\sum_{k=0}^{\infty}\alpha_{k}^{2}<\infty$, then the greedy policy with respect to $Q$ converges to the optimal policy with probability 1 as $k\to\infty$.
@@
Standard Q-learning convergence requires that every state-action pair is updated infinitely often. Our exploring policy $\pi^{\phi}$ prunes actions that are guaranteed to be suboptimal by (7), meaning not every state-action pair is updated. However, the key insight is that for every state $x$, every action that could potentially be optimal is still updated infinitely often (via exploring starts). Thus, while $Q(x,u)$ may not converge to $Q^{*}(x,u)$ for pruned actions, the greedy policy $\pi(x)\in\arg\max_{u\in\mathcal{U}}Q(x,u)$ converges to the optimal policy $\pi^{*}(x)$ with probability 1, since the Q-values of all candidate optimal actions either converge to their true values or are dominated by actions that do. ∎
@@
III-B1 Choice of Weighting $\beta$